Skip to end of metadata
Go to start of metadata

Date

Attendees

Agenda

    • IPR status
    • follow up on Duo for XSEDE (I sent email to vendor but did not hear back yet)
    • Update on central log server XES participants
    • Plan for upgrading XES services at our XES infrastructure provider's data centers (NICS, NCSA, PSC, TACC)
    • Mgr. Coordination 

 

Discussion items

TimeItemWhoNotes
2minIPR StatusScott

Scott provided reminder that tomorrow is end of quarter. 

  • Scott sent templates last week.
  • Get templates back to Scott by Friday.
2minStatus of DuoVictor

Current license ends 4/29

  • Victor has communicated with Duo and is waiting on the pricing information
2minUpdate on central log server XES participantsVictorGary reports we are at 60% XES participation and still adding.
15minPlan for upgrading XES services at our XES infrastructure provider's data centers (NICS, NCSA, PSC, TACC)Victor

Had a first discussion with Ops Mgrs about plan for VM infrastructure to be multiple infrastructure provider data centers than can move VMs between the centers. 

  • Greg had a meeting with Victor, Gary and Tabitha to start the discussion about this last Friday.  
  • Started the discussion with Ops Mgrs and need (google) document to describe the plan and what is needed/wanted.   
  • Mentioned to Tim that we probably need networking space between the centers. 
  • NICS will start this off with next VM infrastructure purchase.   
  • Need to continue discussions about this and have Greg on them.
2minMgr Coordination 

Security and SysOps collaborative meeting every week immediately after Ops Mgmt meeting

XOC back to funding level from before.  Mike mentioned things are operating well and he had one person retire and is working on replacement. Offboarding form was submitted and PM&R team executed the process.

Action items

  •